OpenStudy (anonymous):

A businessman bought a dealership that is incurring a loss of $500,000 a year. He decided to strategize in order to turn the business around. In addition to the $500,000 annual loss, his fixed cost for running the dealership on a monthly basis is $5,000. The number of cars sold per week and their probabilities mimic the outcomes of three coins being flipped. The number of cars sold in a week was observed to be the same as the number of tails that appear in a three coin flip. See the distribution: Number of Tails 0 |1 |2 |3 probablity 1/8 | 3/8 | 3/8 | 1/8

OpenStudy (anonymous):

What is the expected number of cars sold in one week? A. 0 B. 1 C. 1.5 D. 2 E. 2.5

OpenStudy (anonymous):

@amistre64

OpenStudy (whpalmer4):

okay, you are constructing a weighted average here. flipping 0 tails is the same as selling 0 cars, and happens with probability 1/8 flipping 1 tail is the same as selling 1 car, and happens with probability 3/8 flipping 2 tails is the same as selling 2 cars, and happens with probability 3/8 flipping 3 tails is the same as selling 3 cars, and happens with probability 1/8

OpenStudy (whpalmer4):

on average, 1/8 of the time no cars will be sold. 3/8 of the time, 1 car will be sold. 3/8 of the time 2 cars will be sold. 1/8 of the time, 3 cards will be sold.

OpenStudy (whpalmer4):

multiply each number of cars sold by its probability, and add them up.

OpenStudy (whpalmer4):

that's the expected value of cars sold in a week.

OpenStudy (whpalmer4):

thinking of it a slightly different way, if you had 8 weeks, 1 week you sell 0 cars. 3 weeks you sell 1 car. 3 weeks you sell 2 cars. 1 week you sell 3 cars. how many cars are sold over those 8 weeks? what is the average number of cars sold per week?

OpenStudy (anonymous):

i got 2.25

OpenStudy (whpalmer4):

try again.

OpenStudy (whpalmer4):

show your work

OpenStudy (anonymous):

.375 + 1.125+ .75= 2.25

OpenStudy (whpalmer4):

show me how you got those numbers.

OpenStudy (anonymous):

1* 3/8= .375, 2*3/8, 3*1/8

OpenStudy (whpalmer4):

I agree that 1*3/8 = 0.375. I don't agree about the others. in any case, there's no need to convert to decimals at this point in the problem, and it's causing you to make mistakes. just do it as fractions.

OpenStudy (anonymous):

i get the same things

OpenStudy (whpalmer4):

show me your work.

OpenStudy (anonymous):

2 *3 divided by 8, etc.

OpenStudy (anonymous):

i dont see what im doing wrong

OpenStudy (whpalmer4):

look, you aren't doing it correctly, why are you resistant to showing me exactly what you did so I can correct you?

OpenStudy (whpalmer4):

yes, you don't, but I will!

OpenStudy (anonymous):

thats what i did though

OpenStudy (whpalmer4):

let's go through it again: 0 cars are sold 1/8 of the time 1 car is sold 3/8 of the time 2 cars are sold 3/8 of the time 3 cars are sold 1/8 of the time \[0*\frac{1}{8} + 1*\frac{3}{8} + 2*\frac{3}{8} + 3*\frac{1}{8} =\]

OpenStudy (whpalmer4):

what does that equal?

OpenStudy (anonymous):

1.5

OpenStudy (anonymous):

oh i know what i did wrong

OpenStudy (whpalmer4):

yes, 1.5 is correct.

OpenStudy (whpalmer4):

3/8 + 6/8 + 3/8 = 12/8 = 3/2 = 1.5.